In some aquatic ecosystems, nitrate (NO3–) is converted to nitrite (NO2–), which then decomposes to nitrogen and water. As an example of this second reaction, consider the decomposition of ammonium nitrite:
NH4NO2(aq) = N2(g) + 2H2O(l)
What would be the change in pressure in a sealed 10.0 L vessel due to the formation of N2 gas when the ammonium nitrite in 1.40 L of 0.900 M NH4NO2 decomposes at 25.0°C? answer in atm.
1
Expert's answer
2019-12-09T08:04:42-0500

nRT=PV (Mendeleev-Klapeurones law )

n - number of moles, C - concentration M (moles/Litres), V - Volume gas(L), R - constant , P - pressure, T- absolutle Temperature (*K)

n=C * V

n=1.26 moles of NH4NO2

reaction, consider the decomposition of ammonium nitrite:

NH4NO2(aq) = N2(g) + 2H2O(l) 

nNH4NO2=nN2

Product of gas-fase only N2

At 25*C=298*K,

P=nRT/V

P=1.26*298*R/10.0, R=0.082 atm*Litres/moles*K

P=3.079 atm